OpenStudy (anonymous):

Find the solution in the form y=f(x) for the differential equation: dy/dx= 2xy Help is greatly appreciated!

OpenStudy (anonymous):

so first I figure I should separate and integrate: dy/y= 2x dx Int. dy 1/y= Int. 2x dx

OpenStudy (anonymous):

I got: ln(y) +C = x^2 +C2 which is ln(y)= x^2 +C I'm not entirely sure how to get rid of the ln...

OpenStudy (phi):

you can make each side the exponent of e \[ e^{\ln(y)} = e^{x^2+C} \]

OpenStudy (phi):

people rewrite the right side as \[ e^{x^2} e^C = B e^{x^2} \] where B represents the arbitrary constant e^C

OpenStudy (phi):

e^ln(y) simplifies to y

OpenStudy (anonymous):

so y= e^(x^2+C) but I could rewrite as y= ke^(x^2) (using 'k' as your "B")

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you very much for your help! I understand this now, the problem was that it is multiple choice and the answers didn't match up. but I see that I can write y= e^(x^2 +C) as y= ke^(x^2) so it works out!